a deer goes from point a to point b in 7.00 seconds with constant acceleration, traveling 70.0m. Its speed passing through point B is 15.0m / s. The speed at point A was:

A. 35,0 m/s

B. 230 m/s

C. 5,00 m/s

D. 4,29 m/s

E. 260 m/s


1
Expert's answer
2021-01-20T15:49:33-0500

We can find the initial velocity of the deer at point A from the kinematic equation:


"d=\dfrac{v_i+v_f}{2}t,""v_i=\dfrac{2d}{t}-v_f,""v_i=\dfrac{2\cdot 70\ m}{7\ s}-15\ \dfrac{m}{s}=5\ \dfrac{m}{s}."

Answer:

C) "5.0\ \dfrac{m}{s}."